# Disentangled Representation Learning through Geometry Preservation with the Gromov-Monge Gap

### v1

@@ A.3 Conditionally Positive Kernels
Birkhoff’s theorem states that the extremal points of $U_{n}$ are equal to the permutation matrices $P_{n}$. Moreover, a seminal theorem of linear programming [6, Theorem 2.7] states that the minimum of a linear objective on a bounded polytope, if finite, is reached at an extremal point of the polyhedron. Therefore, as $\mathbf{P}^{\star}$ solves Eq. (11), it is an extremal point of $U_{n}$, so it can always be chosen as a permutation matrix. Therefore, the equivalence between GW and GM follows. ∎
To conclude the proof of Prop. 3.4, we simply remark that:
  • $r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and $T\sharp r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$ are uniform, empirical distribution, and supported on the same number of points;
  • The costs of interests [i] or [ii] are either conditionally positive, or their opposite is, as detailed below  Def (A.1).
See 3.5
@@
We first note that the empirical estimator of the distortion is consistent, as both costs [i] or [ii] are continuous, and $\mathcal{X}$ is compact. We then need to study, in both cases, the convergence of $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r_{n},T\sharp r_{n})$ to $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(r_{n},T\sharp r)$.
To that end, we first remark that as, almost surely, $r_{n}\to r$ in distribution, one also has that, almost surely, $T\sharp r_{n}\to T\sharp r$ in distribution. Indeed, since $\mathcal{Y}$ is compact, $T$ is bounded so for any bounded and continuous $f:\mathcal{Y}\rightarrow\mathbb{R}$ and $X\sim r$, $f\circ T(X)$ is well defined and bounded so integrable. Afterwards, one can simply adapt the proof of the almost sure weak convergence of empirical measure based on the strong law of large numbers to show that, almost surely, $T\sharp{r}_{n}\rightarrow T\sharp r$ in distribution. See for instance [39, Theorem 10.4.1].
[i] We start by the (scaled) squared euclidean distances. Up to replacing $r$ by $\alpha^{2}\sharp r$ and $T$ by $T\circ(\tfrac{1}{\alpha^{2}})$, and similarly for $r_{n}$, we can assume without loss of generality that $\alpha=1$. As, almost surely, both $r_{n}\to r$ and $T\sharp{r}_{n}\to T\sharp r$ in distribution, the results follows from [48, Thm 5.1, (e)].
[ii] We continue with the cosine similarity. To that end, we first consider the inner product, i.e., $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, and show that if $p_{n}\to p$ and $q_{n}\to q$ in distribution, then $\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p_{n},q_{n})\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)$. As noticed by Rioux et al. [58, Lemma 2]–in the first version of the paper– the GW for inner product costs can be reformulated as:
@@
    where we remind that that $\mathrm{W}_{2}^{2}$ is the (squared) Wasserstein distance, solution of Eq. (KP) induced by $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$. By virtue of [46, Theorem 2], there exists a constant $C>0$, s.t. we can uniformly bound
@@
As a result, we finally get $\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p_{n},q_{n})\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)$. Finally, we remark that for any $p,q$, $\mathrm{GW}^{\textrm{cos-sim}}(p,q)=\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p,\mathrm{proj}_{S^{d-1}}\sharp q)$, where $\mathrm{proj}_{S^{d-1}}(\mathbf{x})=\mathbf{x}/\|\mathbf{x}\|_{2}$. Using similar arguments invoked previously, as $p_{n}\to p$ in distribution $p$, $\mathrm{proj}_{S^{d-1}}\sharp p_{n}\to\mathrm{proj}_{S^{d-1}}\sharp p$ in distribution, and similarly $\mathrm{proj}_{S^{d-1}}\sharp q_{n}\to\mathrm{proj}_{S^{d-1}}\sharp q$ in distribution. As a result:
$\begin{split}\mathrm{GW}^{\textrm{cos-sim}}(p_{n},q_{n})&=\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p_{n},\mathrm{proj}_{S^{d-1}}\sharp q_{n})\\ &\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(\mathrm{proj}_{S^{d-1}}\sharp p,\mathrm{proj}_{S^{d-1}}\sharp q)\\ &=\mathrm{GW}^{\textrm{cos-sim}}(p,q)\end{split}$ (15)
which yields the desired convergence by using $p_{n}=r_{n}$ and $q_{n}=T\sharp r_{n}$.
Appendix D Proofs of § 3.4
See 3.7
We start by recalling the standard definition of weakly convex function on $\mathbb{R}^{d}$, along with technical lemmas that we will in the proof of Thm. (3.7).
Definition D.1.
@@
Lemma D.2.
@@
We use the fact that a twice continuously differentiable function is convex i.f.f. its hessian is positive semi-definite [8, §(3.1.4)]. Therefore, $f_{\mathbf{A}}$ is convex i.f.f. $\nabla^{2}f_{\mathbf{A}}=\mathbf{A}\geq 0$. If $\lambda_{\min}(\mathbf{A})\geq 0$, then $\mathbf{A}\geq 0$ so $f_{\mathbf{A}}$ is convex, i.e. 00-weakly convex. Otherwise, $f_{\mathbf{A}}-\tfrac{1}{2}\lambda_{\min}(\mathbf{A})\|\cdot\|_{2}^{2}$ has hessian $A-\lambda_{\min}(\mathbf{A})\geq 0$, so it is convex, which yields that $f_{\mathbf{A}}$ is $-\lambda_{\min}(\mathbf{A})$-weakly convex. ∎
Lemma D.3.
@@
As the $f_{i}$ are $\gamma$-weakly convex, $f_{i}+\tfrac{1}{2}\gamma$ is convex, so $\mathbf{x}\mapsto\sup_{i\in I}f_{i}(\mathbf{x})+\tfrac{1}{2}\gamma\|x\|_{2}^{2}=(\sup_{i\in I}f_{i}(\mathbf{x}))+\tfrac{1}{2}\gamma\|x\|_{2}^{2}$ is convex [8, Eq. (3.7)]. Therefore, the $\gamma$-weak convexity of $f$ follows ∎
@@
$\begin{split}\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
As ${r}_{n}$ and $T\sharp{r}_{n}$ are uniform empirical supported on the same number of points, using Prop. C.1, we can reformulate the RHS with permutation matrices, which yields
$\begin{split}\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
From this expression, $\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ can be reformulated as a matrix input function. Indeed, it only depends on the map $T$ via its values on the support of ${r}_{n}$, namely $\mathbf{x}_{1},...,\mathbf{x}_{n}$. Therefore, we write $\mathbf{t}_{i}\vcentcolon=T(\mathbf{x}_{i})$, and define $\mathbf{X},\mathbf{T}\in\mathbb{R}^{n\times d}$ which contain observations $\mathbf{x}_{i}$ and $\mathbf{t}_{i}$ respectively, stored as rows. Then, studying $\mathcal{GM}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ remains to study
@@
where we define $\mathbf{A}_{\mathbf{X},\mathbf{P}}\vcentcolon=\tfrac{1}{n^{2}}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P}-\mathbf{X}\mathbf{X}^{\top})\in\mathbb{R}^{n\times n}$. To study the convexity of this matrix input function, we vectorize it. From [53, Eq. (520)], we note that, for any $\mathbf{M}\in\mathbb{R}^{n\times n}$
@@
To study the convexity of $r$, we study the convexity of each $r_{\mathbf{A}_{\mathbf{X},\mathbf{P}}}(\mathbf{T})\vcentcolon=\mathbf{vec}(\mathbf{T})^{\top}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})\mathbf{vec}(\mathbf{T})$, which are quadratic forms induced by the $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$. This remains to study the (semi-) positive definiteness of the matrices $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$. As each $\mathbf{A}_{\mathbf{X},\mathbf{P}}\in\mathbb{R}^{n\times n}$ is symmetric and square, $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$ is also symmetric and from [53, Eq. (519)] its eigenvalues are the outer products of the eigenvalues of $\mathbf{A}_{\mathbf{X},\mathbf{P}}$ and $I_{n}$, namely
@@
We then remind that each $r_{\mathbf{A}_{\mathbf{X},\mathbf{P}}}$ is the quadratic form defined by $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$, so by applying Prop. D.2, it is $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$-weakly convex, and hence $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex. Therefore, applying Prop. (D.3), $r$ is $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex, in $\mathbb{R}^{d}$. Reminding that $\gamma_{\mathrm{inner}}=\tfrac{1}{n}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$, $r$ is $\tfrac{1}{n}\gamma_{\mathrm{inner}}$ weakly convex. This implies that $\mathbf{T}\mapsto f(\mathbf{T})+\tfrac{1}{n}\gamma_{\mathrm{inner}}\|\mathbf{T}\|_{2}^{2}$ is convex. By reminding that $\mathbf{T}$ stores the $T(\mathbf{x}_{i})$ as rows, $\tfrac{1}{n}\|\mathbf{T}\|_{2}^{2}=\|T\|_{L_{2}(r_{n})}$. Consequently, $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is $\gamma_{\mathrm{inner}}$ in $L_{2}(r_{n})$.
We then study the convexity of $\mathcal{GM}_{r_{n}}^{2}$. We follow exactly the same approach. One has:
$\begin{split}\mathcal{GM}^{2}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
Similarly, studying the convexity of $\mathcal{GM}^{2}_{{r}_{n}}(T)$ remains to study the convexity of the matrix input function:
@@
For both terms, we apply again [53, Eq. (519)]. For the LHS, one has:
@@
Asymptotic. For any $T$, we note that, almost surely, $\|T\|^{2}_{L_{2}(r_{n})}\to\|T\|^{2}_{L_{2}(r)}$. As a result, since convexity is preserved under pointwise convergence and by virtue of Prop. (3.5), we study the (almost sure) convergence of $\gamma_{\textrm{inner},n}$ and $\gamma_{2,n}$.

### v2

@@ B.3 Positivity of the Entropic GMG estimator
Birkhoff’s theorem states that the extremal points of $U_{n}$ are the permutation matrices $P_{n}$. Moreover, a seminal theorem of linear programming (Bertsimas and Tsitsiklis, 1997, Theorem 2.7) states that the minimum of a linear objective on a bounded polytope, if finite, is reached at an extremal point of the polyhedron. Therefore, as $\mathbf{P}^{\star}$ solves Eq. (18), it is an extremal point of $U_{n}$, so it can always be chosen as a permutation matrix. Therefore, the equivalence between GW and GM follows.
To conclude the proof of Prop. B.1, we simply remark that $r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and $T\sharp r_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{T(\mathbf{x}_{i})}$ are uniform, empirical distribution, and supported on the same number of points.
Consistency of the empirical GMG.
We continue by proving a consistency result for the empirical GMG, which we will later use to deduce the asymptotic weak convexity constant from the finite-sample case.
Proposition B.3.
For both $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$ and $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, one has $\mathrm{GMG}_{{r}_{n}}(T)\rightarrow\mathrm{GMG}_{r}(T)$ almost surely.
@@
We first note that the empirical estimator of the distortion is consistent, as both costs are continuous, and $\mathcal{X}$ is compact. We then need to study, in both cases, the convergence of $\mathrm{GW}(r_{n},T\sharp r_{n})$ to $\mathrm{GW}(r_{n},T\sharp r)$.
To that end, we first remark that as, almost surely, $r_{n}\to r$ in distribution, one also has that, almost surely, $T\sharp r_{n}\to T\sharp r$ in distribution. Indeed, since $\mathcal{Y}$ is compact, $T$ is bounded so for any bounded and continuous $f:\mathcal{Y}\rightarrow\mathbb{R}$ and $X\sim r$, $f\circ T(X)$ is well defined and bounded so integrable. Afterwards, one can simply adapt the proof of the almost sure weak convergence of empirical measure based on the strong law of large numbers to show that, almost surely, $T\sharp{r}_{n}\rightarrow T\sharp r$ in distribution. See for instance (Le Gall, , Theorem 10.4.1).
We start with the squared Euclidean distance. As, almost surely, both $r_{n}\to r$ and $T\sharp{r}_{n}\to T\sharp r$ in distribution, the results follows from (Mémoli, 2011, Thm 5.1, (e)).
We continue with the inner product. As noticed by Rioux et al. (2023, Lemma 2)–in the first version of the paper – the GW for inner product costs can be reformulated as:
@@
    where we remind that that $\mathrm{W}_{2}^{2}$ is the (squared) Wasserstein distance, solution of Eq. (KP) induced by $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$. By virtue of (Manole and Niles-Weed, 2024, Theorem 2), there exists a constant $C>0$, s.t. we can uniformly bound
@@
As a result, we finally get $\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p_{n},q_{n})\to\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)$.
Weak convexity.
Finally, we demonstrate some useful results on weakly convex functions on $\mathbb{R}^{d}$.
Definition B.4.
@@
From the definition, we see that if $f$ is $\gamma$-weakly convex, than $f$ is also $\gamma^{\prime}$ weakly convex for any $\gamma^{\prime}\geq\gamma$. This naturally extends to weakly convex functionals $\mathcal{F}$ on $L_{2}(r)$.
Lemma B.5.
@@
We use the fact that a twice continuously differentiable function is convex i.f.f. its hessian is positive semi-definite (Boyd and Vandenberghe, 2004, §(3.1.4)). Therefore, $f_{\mathbf{A}}$ is convex i.f.f. $\nabla^{2}f_{\mathbf{A}}=\mathbf{A}\geq 0$. If $\lambda_{\min}(\mathbf{A})\geq 0$, then $\mathbf{A}\geq 0$ so $f_{\mathbf{A}}$ is convex, i.e. 00-weakly convex. Otherwise, $f_{\mathbf{A}}-\tfrac{1}{2}\lambda_{\min}(\mathbf{A})\|\cdot\|_{2}^{2}$ has hessian $A-\lambda_{\min}(\mathbf{A})\geq 0$, so it is convex, which yields that $f_{\mathbf{A}}$ is $-\lambda_{\min}(\mathbf{A})$-weakly convex. ∎
Lemma B.6.
@@
As the $f_{i}$ are $\gamma$-weakly convex, $f_{i}+\tfrac{1}{2}\gamma$ is convex, so $\mathbf{x}\mapsto\sup_{i\in I}f_{i}(\mathbf{x})+\tfrac{1}{2}\gamma\|x\|_{2}^{2}=(\sup_{i\in I}f_{i}(\mathbf{x}))+\tfrac{1}{2}\gamma\|x\|_{2}^{2}$ is convex (Boyd and Vandenberghe, 2004, Eq. (3.7)). Therefore, the $\gamma$-weak convexity of $f$ follows ∎
Let’s now proceed to prove the main Thm. (3.5).
@@
$\begin{split}\mathrm{GMG}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
As ${r}_{n}$ and $T\sharp{r}_{n}$ are uniform empirical supported on the same number of points, using Prop. B.2, we can reformulate the RHS with permutation matrices, which yields
$\begin{split}\mathrm{GMG}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle-\langle T(\mathbf{x}_{i}),T(\mathbf{x}_{j})\rangle|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\langle\mathbf{x}_{i},\mathbf{x}_{i^{\prime}}\rangle-\langle T(\mathbf{x}_{j}),T(\mathbf{x}_{j^{\prime}})\rangle|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
From this expression, $\mathrm{GMG}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ can be reformulated as a matrix input function. Indeed, it only depends on the map $T$ via its values on the support of ${r}_{n}$, namely $\mathbf{x}_{1},...,\mathbf{x}_{n}$. Therefore, we write $\mathbf{t}_{i}\vcentcolon=T(\mathbf{x}_{i})$, and define $\mathbf{X},\mathbf{T}\in\mathbb{R}^{n\times d}$ which contain observations $\mathbf{x}_{i}$ and $\mathbf{t}_{i}$ respectively, stored as rows. Then, studying $\mathrm{GMG}^{\langle\cdot,\cdot\rangle}_{{r}_{n}}$ remains to study
@@
where we define $\mathbf{A}_{\mathbf{X},\mathbf{P}}\vcentcolon=\tfrac{1}{n^{2}}(\mathbf{P}^{\top}\mathbf{X}\mathbf{X}^{\top}\mathbf{P}-\mathbf{X}\mathbf{X}^{\top})\in\mathbb{R}^{n\times n}$. To study the convexity of this matrix input function, we vectorize it. From (Petersen and Pedersen, 2008, Eq. (520)), we note that, for any $\mathbf{M}\in\mathbb{R}^{n\times n}$
@@
To study the convexity of $r$, we study the convexity of each $r_{\mathbf{A}_{\mathbf{X},\mathbf{P}}}(\mathbf{T})\vcentcolon=\mathbf{vec}(\mathbf{T})^{\top}(\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n})\mathbf{vec}(\mathbf{T})$, which are quadratic forms induced by the $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$. This remains to study the (semi-) positive definiteness of the matrices $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$. As each $\mathbf{A}_{\mathbf{X},\mathbf{P}}\in\mathbb{R}^{n\times n}$ is symmetric and square, $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$ is also symmetric and from (Petersen and Pedersen, 2008, Eq. (519)) its eigenvalues are the outer products of the eigenvalues of $\mathbf{A}_{\mathbf{X},\mathbf{P}}$ and $I_{n}$, namely
@@
We then remind that each $r_{\mathbf{A}_{\mathbf{X},\mathbf{P}}}$ is the quadratic form defined by $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$, so by applying Prop. B.5, it is $\mathbf{A}_{\mathbf{X},\mathbf{P}}\otimes I_{n}$-weakly convex, and hence $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex. Therefore, applying Prop. (B.6), $r$ is $\tfrac{1}{n^{2}}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$-weakly convex, in $\mathbb{R}^{d}$. Reminding that $\gamma_{\mathrm{inner}}=\tfrac{1}{n}(\lambda_{\max}(\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top}))$, $r$ is $\tfrac{1}{n}\gamma_{\mathrm{inner}}$ weakly convex. This implies that $\mathbf{T}\mapsto f(\mathbf{T})+\tfrac{1}{n}\gamma_{\mathrm{inner}}\|\mathbf{T}\|_{2}^{2}$ is convex. By reminding that $\mathbf{T}$ stores the $T(\mathbf{x}_{i})$ as rows, $\tfrac{1}{n}\|\mathbf{T}\|_{2}^{2}=\|T\|_{L_{2}(r_{n})}$. Consequently, $\mathrm{GMG}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is $\gamma_{\mathrm{inner}}$ in $L_{2}(r_{n})$.
We then study the convexity of $\mathrm{GMG}_{r_{n}}^{2}$. We follow exactly the same approach. One has:
$\begin{split}\mathrm{GMG}^{2}_{{r}_{n}}(T)&=\tfrac{1}{n^{2}}\sum_{i,j=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}\\ &-\tfrac{1}{n^{2}}\min_{\mathbf{P}\in P_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}\tfrac{1}{2}|\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}^{2}-\|T(\mathbf{x}_{i})-T(\mathbf{x}_{j})\|_{2}^{2}|^{2}|^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}\end{split}$
Similarly, studying the convexity of $\mathrm{GMG}^{2}_{{r}_{n}}(T)$ remains to study the convexity of the matrix input function:
@@
For both terms, we apply again (Petersen and Pedersen, 2008, Eq. (519)). For the LHS, one has:
@@
Asymptotic. For any $T$, we note that, almost surely, $\|T\|^{2}_{L_{2}(r_{n})}\to\|T\|^{2}_{L_{2}(r)}$. As a result, since convexity is preserved under pointwise convergence and by virtue of Prop. (B.3), we study the (almost sure) convergence of $\gamma_{\textrm{inner},n}$ and $\gamma_{2,n}$.